OpenStudy (anonymous):

Solve rational equation:

OpenStudy (anonymous):

\[\frac{ 2x }{ 2x-3 }+ \frac{ 4 }{ x }=\frac{ x-18 }{ x(2x-3) }\]

OpenStudy (anonymous):

@jim_thompson5910 this is my last one and I am out of hear for the night

OpenStudy (anonymous):

First you have to identify the LCD and make the other terms simmilar!

OpenStudy (anonymous):



OpenStudy (anonymous):

(2x)/(2x-3) + 4/x = (x-18)/(x(2x-3)) 2x + (2x-3)(4)/x = (x-18)/x 2x^2 + 8x - 12 = x -18 2x^2 + 7x + 6 = 0 2x^2 + 4x + 3x + 6 = 0 2x(x + 2) + 3(x+2) = 0 (2x + 3)(x + 2) = 0 2x + 3 = 0 and x + 2 = 0 2x = -3 and x = -2 x = -3/2

OpenStudy (anonymous):

@franciscanmonk is the reason you didn't use 2x-3 because -3 is the restriction

OpenStudy (anonymous):

no. i used (2x - 3) it disappeared after line 4 of my solution. i distributed it and then the -18 got transferred and so...the solution goes.

OpenStudy (anonymous):

^ what does that mean

OpenStudy (anonymous):

(2x)/(2x-3) + 4/x = (x-18)/(x(2x-3)) I distribiuted (2x-3) and thus it is gone in the 2nd line.. i 2x-3 is gone from the denominator of the 1st one and the 3rd one. 2x + (2x-3)(4)/x = (x-18)/x I distributed x so that it will vanish from the denominators of the others. 2x^2 + 8x - 12 = x -18 <--- I distributed (2x-3) and got 8x-12 there is no restriction here.